OpenStudy (anonymous):

find the distance between the point (-5,8) and (1,16) whats the distance formula??

OpenStudy (anonymous):

the distance formula is basically the pythagorean theorem

OpenStudy (anonymous):

which is?...

OpenStudy (anonymous):

the square of the length of the hypotenuse of a right triangle is equal to the sum of the squares of the lengths of its legs, or the length of the hypotenuse is equal to \[\sqrt{x^2+y^2}\]

OpenStudy (kira_yamato):

\[\sqrt{(x_1 - x_2)^2 + (y_1 - y_2)^2}\]

OpenStudy (anonymous):

to compute distances between points in the graph, you would apply that theorem. \[Distance = \sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (kira_yamato):

Where the coordinates of the 2 pts are (x1,y1) and (x2,y2)

OpenStudy (anonymous):

thank youuuu

OpenStudy (anonymous):

E=Dy/Dx =(8-16)/(-5-1) =-1.3333